A farmer wanted to increase his crop production and decided to add Potassium-rich fertilizer to his field. 
Rank the following fertilizers based on increasing order of K% (1 - lowest, 5 - highest)
KCl, K2SO4, KNO3, K2SO4.MgSO4, KCl + NaCl + MgSO4?




1
Expert's answer
2022-02-14T06:22:01-0500

Molar mass of :

KCl =74.5.......5(52.3%)

K2SO4=174..…..3(22.4%)

KNO3=101.........4 (38.6%)

K2SO4.MgSO4=294....….2 (13.2%)

KCl+NaCl+MgSO4=306.5.......1(12.73%)